OpenStudy (anonymous):

A particle moves in a straight line so that, t seconds after passing through a fixd point O, its velocity, v m/s, is given by \[v=\frac{60}{(3t+4)^{2}}\] (iii) Find an expression for the displacement of the particle from O, t seconds after it has passed through O. I will post my working, hopefully you'll tell me where I went wrong. A particle moves in a straight line so that, t seconds after passing through a fixd point O, its velocity, v m/s, is given by \[v=\frac{60}{(3t+4)^{2}}\] (iii) Find an expression for the displacement of the particle from O, t seconds after it has passed through O. I will post my working, hopefully you'll tell me where I went wrong. @Mathematics

OpenStudy (anonymous):

\[\int\frac{60}{(3t+4)^2}dx=60\cdot\int(3t+4)^{-2}\]

OpenStudy (anonymous):

\[=60[\frac{(3t+4)^{-1}}{(-1)(3)}]\]

OpenStudy (anonymous):

\[=60[\frac{1}{3t+4}\cdot\frac{1}{-3}]\]

OpenStudy (anonymous):

haha forgot the +C. \[=60[\frac{1}{-9t-12}+C]\]

OpenStudy (anonymous):

\[=\frac{60}{-9t-12}+60C\] which is displacement against time right? \[s=\frac{60}{-9t-12}+60C\]

OpenStudy (anonymous):

what's the answer in the book

OpenStudy (anonymous):

When t=0, s=0 also. \[0=\frac{60}{-9(0)-12}+60C\] \[0=\frac{60}{-12}+60C\] \[60C=\frac{60}{12}\]

OpenStudy (anonymous):

The answer in the book is: \[s=5-\frac{20}{3t+4}\]

OpenStudy (anonymous):

don't multiply 60 by c... c1+c2+c3=c same as 60c will just equal another constant

OpenStudy (anonymous):

so you should get just c=60/12

OpenStudy (anonymous):

thats where the 5 comes then above in your integral if you pull out a 3 from your denominator you should be able to reduce 6 to -20

OpenStudy (anonymous):

continuing my working... \[C=\frac{1}{12}\] So the equation is... \[s=\frac{60}{-9t-12}+60(\frac{1}{12})\] \[s=\frac{60}{-9t-12}+5\]

OpenStudy (anonymous):

Yeah, multiplying the C wasn't necessary, but it doesn't give a wrong answer.

OpenStudy (anonymous):

on the denominator pull out a -3 and see what you get =]

OpenStudy (anonymous):

technically your answer is correct just not fully simplified

OpenStudy (anonymous):

thank your book for not giving all the forms that it could be lol

OpenStudy (anonymous):

\[s=-\frac{20}{3t+4}+5\]

OpenStudy (anonymous):

BALLS That is correct!